what mass of water will be produced from 9.6 g of calcium hydroxide reacting with hcl
1
Expert's answer
2020-05-20T13:43:46-0400

Ca(OH)2 + 2HCl = CaCl2 + 2H2O

9.6 g x g

74 g 2*18 g

x = 9.6 * 2 * 18 / 74 = 4.67 g H2O